help please!
Solve the equation for x and y. x+2y=-5 x=2y-1

choices are: (5,3) (5,0) (-3,-1) (1,-3)

Found 2 solutions by praseenakos@yahoo.com, funmath:
Answer by praseenakos@yahoo.com(507)

QUESTION:
Solve the equation for x and y. x+2y=-5 x=2y-1

choices are: (5,3) (5,0) (-3,-1) (1,-3)
ANSWER:
x+2y=-5 ----------------------(1)
x=2y-1 -----------------------(2)
(5,3)
Input the values x = 5 and y = 3 in both the equations;
(1) ==> 5+ 2 *3 = 5 + 6 = -11, which is not equal to the right side (ie: -5)

If the given point does not satisfy any one of the given equation, then it is not a solution.
Now take the second point (5,0) and substitute in the given equations.

==> 5 + 2 * 0 = 5 which is not equiatl to -5

so (5,0) is not a solution.

Now take the third point (-3, -1)

==> -3 + 2 * -1 = -3 + -2 = -5, which is equal to the right side.
Again substitute these values in the second equation,

==> -3 = 2 * -1 - 1

==> -3 = -2 - 1
==> -3 = -3

That is both sides of the equations are same.

So the point (-3, -1 ) is a solution for the given equations.

We have one more point( 1, -3 ), let's check that also.
(1)==> 1+ 2 * -3 = -5
==> 1 + -6 = -5
==> -5 = -5, that is the point satisfies the first equation.

Now check this point with the second equation.
==> 1 = 2 * -3 - 1

==> 1 = -6 - 1
==> 1 = -7, which is not correct.
so (1, -3) is not a solution.

Conclusion:
The point (-3, -1 ) satisfies both the equation, so it the solutionm for the given pair of equation.
So solution is x = -3 and y = -1







Answer by funmath(2933)

E1: x+2y=-5
E2: x=2y-1
Substitute x=2y-1 into E1 and solve for y.
2y-1+2y=-5
4y-1=-5
4y-1+1=-5+1
4y=-4
4y/4=-4/4
y=-1
Now substitute y=-1 into E2 and solve for x.
x=2(-1)-1
x=-2-1
x=-3
The solution is (-3,-1)
Happy Calculating!!!